OpenStudy (anonymous):

simplify and write expression with out a negative a^3b^2c^-2/ a^5b^2c^4=

hartnn (hartnn):

use this : \(\huge \frac{a^x}{a^y}=a^{x-y}\) so what would be \(\huge \frac{a^3}{a^5} \quad \frac{b^2}{b^2}\quad \frac{c^{-2}}{c^4}=?\)

OpenStudy (anonymous):



hartnn (hartnn):

nopes, \(\huge a^{(3-5)} b^{(2-2)} c^{(-2-4)} = ?\)

OpenStudy (anonymous):



hartnn (hartnn):

3-5 = ?

OpenStudy (anonymous):

-2

hartnn (hartnn):

so it would be \(a^{-2}\) isn't it ? now since u want positive powers, u use \(\huge x^{(-y)}=\frac{1}{x^y}\)

hartnn (hartnn):

so what would be \(\huge a^{-2}c^{-6}= ?\)

OpenStudy (anonymous):

1/a^2c^6

hartnn (hartnn):

yup , thats correct ....

OpenStudy (anonymous):

thanks again!

hartnn (hartnn):

welcome again :)